### v2

# Multi-site Integrated Optical Addressing of Trapped Ions

Abstract
One of the most effective ways to advance the performance of quantum computers and quantum sensors is to increase the number of qubits or quantum resources in the system. A major technical challenge that must be solved to realize this goal for trapped-ion systems is scaling the delivery of optical signals to many individual ions. In this paper we demonstrate an approach employing waveguides and multi-mode interferometer splitters to optically address multiple ${}^{171}\textrm{Yb}^{+}$ions in a surface trap by delivering all wavelengths required for full qubit control. Measurements of hyperfine spectra and Rabi flopping were performed on the E2 clock transition, using integrated waveguides for delivering the light needed for Doppler cooling, state preparation, coherent operations, and detection. We describe the use of splitters to address multiple ions using a single optical input per wavelength and use them to demonstrate simultaneous Rabi flopping on two different transitions occurring at distinct trap sites. This work represents an important step towards the realization of scalable integrated photonics for atomic clocks and trapped-ion quantum information systems.
preprint: APS/123-QED
Ions stored in RF Paul traps Paul:1953 have provided an effective platform for quantum information due to their stability wu:2021 , long coherence times wang:2021 , and high fidelities clark:2021 . In addition to computing srinivas:2021 ; leung:2018 , ion traps have also been used for atomic clocks brewer:2019a ; Burt:2021Nat , quantum networks moehring:2007 ; Nichol:2022Nat , quantum sensors marciniak:2022 , and fundamental science Zhang:2018NC ; Pinkas:2023NP . The development of surface traps chiaverini:2005 ; blain:2021 furthered the potential for advancing these applications by enabling scaling to more ions. Using microfabrication capabilities originally developed by the semiconductor industry, surface ion traps were fabricated that could support many-ion quantum computers pino:2021 ; noel:2022 . These traps also provided a convenient platform for integrating other electronic and optical technologies moody:2022 , which are necessary for the individual addressing and readout of growing numbers of ions. These technologies include photonic waveguides, waveguide splitters Vasques:2022PRL , grating couplers mehta:2016 ; niffenegger:2020 ; ivory:2021 , detectors setzer:2021 ; Reens:2022PRL ; todaro:2021 , and modulators hogle:2023 , all of which combined can support quantum systems with many ions as well as applications which require low size, weight, and power (SWaP).
While many of these proof-of-principle experiments focused on the operation of a single ion site, simultaneous independent control of multiple ions at several trap sites had until now not been realized. Here, we demonstrate a room temperature ion trap and simultaneous optical addressing of three ${}^{171}\textrm{Yb}^{+}$ions in independent wells, using only light delivered through waveguides. The light addressing two of the three ions comes from a single split input, forming a multi-ion ensemble. Multi-mode interferometer (MMI) splitters are employed to equally divide the light from the input waveguide into two separate waveguides, which are routed to output couplers individually addressing the trapped ions. Other splitting ratios are possible by varying the interaction length in the MMI splitters, as are other fanout ratios (e.g. 1$\times$×N). This technique could be employed to deliver light to a much larger number of ions, ultimately limited by the power handling capacity of waveguides. A demonstration of 300 mW of 729 nm light in a silicon nitride waveguide mehta:2020 shows that many sites could be supported from a single launch. Large collections of ions are particularly interesting for atomic clocks, where it has been shown that separately interrogated ensembles of ions can achieve sensitivities that scale as $(\alpha N)^{-m/2}$, where $N$ is the number of ions in each of $m$ ensembles, and $\alpha$ is a protocol-dependent constant borregaard:2013 .
In this experiment we performed hyperfine spectroscopy and Rabi flopping on the E2 clock transition, consisting of Doppler cooling, state preparation, coherent operations, and detection. Only the photoionization laser beams were delivered via free space. This required multiple wavelengths ranging from UV (369 nm) to NIR (935 nm). We demonstrated the bluest wavelength used for integrated optical addressing to date, overcoming challenges related to the lithography of the grating couplers and propagation losses of the waveguides. Additionally, we measured simultaneous Rabi flopping for two different transitions occurring at different sites, which is a key capability for quantum systems that require magnetic field calibration. We also investigated optical cross-talk between sites.
I Experimental setup
Figure 1: Experimental setup. a, Conceptual illustration of a surface trap with ions and integrated beams. b, SEM image of the trapping region of the micro-fabricated surface trap, consisting of multi-color waveguide output couplers (oval structures within RF rails), electrodes (green) and RF rails (blue). Single photon avalanche diodes (SPADs) are located at the center of trapping sites and are shielded by a grounded metal mesh to protect the ion from electric field perturbations when the SPADs are operational (see Methods). c, SEM image for an MMI splitter that routes light to waveguide output couplers (indicated in blue & dark blue). Similar splitters are also applied for other colors. d, SEM image of grating output couplers. e, ${}^{171}\textrm{Yb}^{+}$energy level diagram, indicating the three different integrated wavelengths used in the experiment.
Fig. 1b shows a false-color scanning electron microscope (SEM) image of multiple sites in the microfabricated surface trap. Each site was optically addressed by light from four output couplers, allowing for the necessary state manipulation needed for a clock based on ${}^{171}\textrm{Yb}^{+}$. An energy level diagram of the relevant levels is shown in Fig. 1e.
Ions were trapped 50 $\mu$m above the surface with a 200 $\mu$m spacing between trap sites. Light was coupled into the waveguides through input gratings located far from the trapping region and routed to the output gratings. Single-layer aluminium oxide waveguides were used for the UV light at 369 nm and 435 nm, while silicon nitride waveguides were used for 760 nm and 935 nm. These materials were chosen for their low-loss at these wavelengths and for ease of fabrication. Light from three output couplers converged at each of three trap sites along the trap axis, allowing for three ions to be cooled and manipulated using only light from the integrated photonics.
[…]
Ions were trapped at room temperature with a vacuum pressure of $\approx$1×$\times 10^{-11}$ Torr. A magnetic field of 5.5 G ($1.4\hat{\textrm{{x}}}+0.7\hat{\textrm{{y}}}+5.3\hat{\textrm{{z}}}$) was applied to optimize both ion fluorescence and coupling to the quadrupole transition. This bias magnetic field also broke the degeneracy of the Zeeman sublevels of the excited states.
A defining feature of these integrated photonics is that the waveguides for two of the three trap sites were fed via a common input. For each wavelength, an integrated splitter, as shown in Fig. 1c, was used to equally divide the input light into two waveguide channels which were routed to the output gratings of the two sites. The MMI splitters were measured to have low optical losses of $<0.2$ dB per splitter. Integrated splitters allow for the control of numerous ions at different sites sites via a single input, providing a path for supporting many-ion ensembles with only a few optical inputs.
II Multi-ion trapping and waveguide characterization
State detection is particularly challenging in surface traps using integrated waveguides because light emitted from output couplers can scatter into the imaging system and overwhelm the ion fluorescence. In a standard detection scheme Streed:2011PRL ; niffenegger:2020 , ions fluoresce at the same wavelength as the detection laser, rendering wavelength filters too broad to distinguish between the two. The light output from a grating is highly polarized and therefore polarization filtering is possible, however it was found to be insufficiently discriminating in this experiment.
In addition to a conventional imaging system with both a CCD camera (Andor Zyla) and photo-multiplier tube (PMT, Hamamatsu H10682-210) for collecting fluorescence normal to the trap surface, a side-imaging lens was added to collect light along the $y$-axis. This geometry avoids collecting the off-axis scattered photons from the cooling/detection beams that emerge at $45^{\circ}$ from the waveguide output couplers, which otherwise overwhelms the ion fluorescence signal when imaged from above. As a standard 6-inch octagonal vacuum chamber was used, the distance from the ion to the objective placed at a side viewport was larger by a factor of 5 compared to conventional overhead imaging, and therefore was much less efficient at collecting photons due the reduced numerical aperture (see Methods). With the conventional overhead imaging system and free-space light delivery, the total count rate on the PMT was 30 kcps (kilocounts per second), with negligible background noise ($<$0.1 kcps). For waveguide-delivered light and overhead imaging, the signal-to-noise ratio was $\ll$1. Using side-imaging and waveguide-delivered light, the total count rate was 5-8 kcps with high but manageable background noise ($\sim$2 kcps) from scattered photons.
Figure 2: Multi-site optical addressing and integrated beam profiles. a, Image of multiple trapped ions taken with a CCD camera from the side of the trap. The top image shows three ions illuminated by light from waveguides (i, ii, and iii), with the left two (i and ii) using light split from single sources corresponding to the different required wavelengths. The fourth ion highlighted with the dashed red circle was illuminated with free-space beams at the loading site. b, Beam profiles for the waveguide-delivered 369 nm and 435 nm beams. The intensities were calculated using the ion fluorescence (369 nm beam) and Rabi frequency (435 nm beam). A 3D contour plot shows normalized intensity along the x-z plane. The $x$-axis origin is at the center of the middle SPAD, consistent with the coordinates in part a. The beams intersected at 50 $\mu$m from the surface and near the center of the SPAD. The white-dashed line represents RF null position on z-axis. c, Measured profiles along the trap axis of the 369 nm and 435 nm at the height corresponding to the optimized interaction strength (used in Fig. 3). Error bars show the standard error of the mean. d, Simulated waveguide profiles for Doppler (left) and clock transition (right) beams. The $x^{\prime}$-axis in this figure is along the line between waveguide gratings. For the $z$-axis, 0 is the trap surface.
With the side-imaging system, the objective lens collected and focused light onto either a CCD camera (Andor Luca) or a 32 channel, linear PMT array (Hamamatsu H11659-200) with a multi-pinhole filter for the three waveguide sites. The 100 $\mu$m  pinholes spatially filtered the ion fluorescence from background scattering to improve the signal-to-noise ratio (SNR) on the PMTs. Ultimately this arrangement achieved SNRs between 5 and 10 while supporting detection times on the order of milliseconds (4 ms for the measurement in Fig. 3). Fig. 2a shows fluorescence images of the ions illuminated by both free-space and waveguide beams. Each waveguide-illuminated ion, labeled (i) to (iii), was separated by a distance of 200 $\mu$m and placed in a local potential minimum created by the surrounding electrodes. To trap multiple ions simultaneously, we loaded each ion sequentially. The first ion was loaded at the loading hole position (marked with a red circle in Fig. 2a) using free-space beams (Doppler and repump) with a single-site potential solution (see Methods). This loaded ion was then shuttled towards the far-right position (iii) by applying the appropriate voltage waveforms to the interior control electrodes. This shuttling process was carried out without the use of cooling beams between sites, but once the ion reached the target position it was cooled by light from the output gratings. We subsequently loaded new ions by applying an additional potential minimum and repeating this process until three ions were stored at locations (i) to (iii).
While the Rabi flopping measurements in this paper used PMTs, we also fabricated SPADs for each ion-trapping site for future integrated detection. We shuttled and then detected ions with the SPADs using free space lasers, but ultimately could not operate the waveguides and SPADs simultaneously due to scattering from the waveguides that dominated the count rate.
The applied control and RF voltages produced an axial frequency of $2\pi\times 1.02$ MHz and radial frequency of $2\pi\times 3.52$ MHz, corresponding to a radial trap depth of 71 meV. Positioning the DC electrodes inside the RF rails led to a significant increase in DC electrode efficacy, such that static voltages within ±$\pm 1$V were sufficient for achieving a 1 MHz axial frequency.
The 369 nm and 935 nm beams were used for cooling, state preparation, and detection of the ions, while the 435 nm beam was used for performing coherent operations on the E2 transition. Additional beams required for ion loading (399 nm and 393 nm) were delivered via free-space optics. This was done to reduce trap complexity, but future devices could easily incorporate waveguides that support photoionization. Waveguides for delivering 760 nm repump light from the fourth output coupler were fabricated but never used. This wavelength can be used for depopulating the F state when a background gas collision transitions the ion to this state, but it was not necessary for these experiments given the several-hour lifetime achieved at the ultra-high vacuum pressures in the chamber.
All waveguide output gratings were designed to emit beams that overlapped at the trapping site 50 $\mu$m above the surface. While the 369 nm and 435 nm gratings were designed to focus at the ion’s location, the repump beams were intentionally unfocused in order to ensure they overlapped with the ion position. Fig. 2b shows the measured waveguide profiles for 369 nm and 435 nm in $\hat{\textrm{{x}}}$ and $\hat{\textrm{{z}}}$. The height of the ion (z) is controlled by adding an electric shim field in $\hat{\textrm{{z}}}$, with the resulting distances measured using the side-imaging camera.
As shown in Fig. 2c, the two beam profiles have closely overlapping foci near the 50 $\mu$m ion height, and near the center of the integrated SPADs. The 369 nm integrated beam profile in Fig. 2c is measured using ion fluorescence well below saturation intensity, while ensuring that the signal is not cut off by the side-imaging pinholes. For the 435 nm quadrupole beam, the beam profile was characterized by measuring the Rabi frequency as a function of ion position. In both cases the profiles were fit to a Gaussian, showing a full-width-half-maximum beam width of 5.26 $\mu$m and 5.25 $\mu$m respectively.
The in-situ measurements of the integrated beams agreed well with independent microscope measurements. Fig. 2d shows the theoretically simulated waveguide output coupler profile for 369 nm and 435 nm beams, where the beams are overlapping around 50 $\mu$m as designed. The $x^{\prime}$-axis in this figure is not the original $x$-axis but instead corresponds to the projection of the $x$-axis on a plane that is perpendicular to the trap surface and runs through the centers of the gratings.
Figure 3: Coherent qubit operations. a, Schematic of split and single waveguide output grating couplers. Blue output grating couplers are routed from the splitter and share the same input, whereas orange is independent. b, Simultaneous Rabi flopping measurements of the three waveguide-illuminated ions. The trace colors correspond to the similarly colored gratings in a. The average ion population of the D-state ($m_{\textrm{F}}=-1$) was measured while scanning the 435 nm pulse time. The Rabi flops were fit with a single-ion decoherence Rabi oscillation formula in the Lamb-Dicke regime Semenin:2022OLP (see Methods). c, Same as b but with the 435 nm light tuned to the ($m_{\textrm{F}}=0$) state for the two ion ensemble (light and dark blue), while the 435 nm light for the single ion (orange) remained tuned to the $m_{\textrm{F}}=-1$ state. Each data point was the average of three trials of 200 measurements of state-detection for each trace. Error bars show the standard error of the mean.
III Simultaneous multi-ion/multi-state addressing
In these experiments, ensembles consisting of 2$+$1 trapped ions (2 ions with shared waveguide inputs and a single ion with independent control) were probed simultaneously. Each ion was first optically pumped to |$\ket{2\textrm{S}_{1/2},F=0}$ and then a 435 nm laser resonant with the quadrupole transition was used to demonstrate Rabi flopping with |$\ket{2\textrm{D}_{3/2},F=2}$ (Fig. 3). The ion positions for the split beams (blue traces) were set to maximize the interaction strength for each ion within each 435 nm beam. Ideally, this would correspond to the maximum intensity, but other factors like micromotion could slightly offset these optima. The two ions exhibited similar Rabi rates ($\delta\Omega=0.064(4)\Omega$), indicating a near 50/50 split for the MMI. Though not necessary in this case, the interaction strengths could have been fine-tuned by adjusting the position of each ion to equalize their Rabi rates. The orange data indicates the independently addressed ion. Each data set was fit with a single-ion, two-state transition model that accounted for the average motional quanta $\bar{n}$ of the ion in the Lamb-Dicke regime.
In practical realizations of quantum computers, quantum sensors, or atomic clocks, it may be beneficial to probe multiple transitions with different sensitivities in order to calibrate magnetic fields. To demonstrate this capability, we simultaneously probed two different transitions (one with each ensemble) as shown in Fig. 3c. This measurement was similar to Fig. 3b, but with a different optical qubit frequency used for the ions addressed by the split waveguides. These ions (dark and light blue) transitioned from |$\ket{2\textrm{S}_{1/2},F=0,m_{\textrm{F}}=0}$ to |$\ket{2\textrm{D}_{3/2},F=2,m_{\textrm{F}}=0}$ (the ${}^{171}\textrm{Yb}^{+}$clock transition), while the single-waveguide ion transitioned from |$\ket{2\textrm{S}_{1/2},F=0,m_{\textrm{F}}=0}$ to |$\ket{2\textrm{D}_{3/2},F=2,m_{\textrm{F}}=-1}$. Due to their different coupling coefficients, they exhibited different Rabi rates. The frequency required for the $\delta m=0$ transition state is first-order insensitive to magnetic fields and is naturally well-overlapped (the difference in Rabi rates is $\delta\Omega=0.039(4)\Omega$). This further confirmed the near 50/50 power split performance of the MMI splitter, as the Rabi rates became closer when the transition was not susceptible to spatial magnetic field gradients.
The ions were not cooled to the motional ground state in this experiment ($\bar{n}\approx 30$, based on previous Rabi fits), which leads to decay in the Rabi contrast as pulse time increases wineland:1998 . This $\bar{n}$ is higher than the Doppler limit ($\approx$10 quanta), primarily due to insufficient cooling times (discovered later in the experiment) and less optimal Doppler cooling caused by the fixed polarization and $k$-vector of the light emitted from the output coupler relative to the applied magnetic field, based on comparisons to the lower $\bar{n}$ measured with free space beams. Heating rates were measured at site (ii) to be 1-3 q/ms (see Methods). The detection fidelity achieved in this experiment is lower than typical, primarily due to collection efficiency limitations of the side-imaging system and the higher $\bar{n}$. Our chamber setup unavoidably increased the distance between the side-imaging objective lens and the trapped ion, making it challenging to use high numerical aperture objective lenses. Additionally, while greatly improved by using side-imaging, there remained some scattering from the waveguide output couplers that contributed to background noise.
IV Cross-talk characterization
Figure 4: Optical cross-talk a, Measured cross-talk between the single ion ensemble (orange) and two ion ensemble (light and dark blue). Rabi oscillations with the D-state ($m_{F}=-1$) were measured while addressing the single ion ensemble (orange). Cross-talk was evident in the slow transition rate of the nearer ion (light blue) in the two ion ensemble (light and dark blue), and was more attenuated for the more distant ion. b, A similar measurement as a, but instead addressing the two ion ensemble (light blue and dark blue). Cross-talk was evident in the slow transition rate of the single ion ensemble (orange). c, Relative cross-talk intensities of transition light. Ratios were calculated from fits of Rabi frequencies and plotted for each trap site.
An important consideration in any multi-ion system is optical cross-talk, as the unintended light on neighboring ions introduces errors in the clock or qubit. Cross-talk errors for integrated addressing of a single ion in a chain have been measured as a function of displacement from the center of the beam profile mehta:2016 , as well as in a multi-zone trap mehta:2020 .
To measure cross-talk, light was applied at one site while Rabi transitions were measured at all sites. The ion(s) of a single ensemble were probed using the |$\ket{2\textrm{S}_{1/2},F=0,m_{\textrm{F}}=0}$ to |$\ket{2\textrm{D}_{3/2},F=2,m_{\textrm{F}}=-1}$ transition, in the same way as Fig. 3b, while the corresponding optical cross-talk on the ion(s) of the other ensemble was observed. Fig. 4a shows that addressing the single ion ensemble resulted in a Rabi frequency $\Omega_{0}$ for the single ion and a cross-talk-induced Rabi frequency $\Omega_{c}\sim 0.05\Omega_{0}$ on the closest ion of the other ensemble. This corresponds to a relative cross-talk intensity of $I_{c}/I_{0}=0.0026(1)$, where $I_{0}\propto\Omega_{0}^{2}$ is the driven intensity and $I_{c}$ is the cross-talk induced intensity.
It is notable that only ion (ii) of the ensemble showed measurable optical cross-talk. If scattering due to waveguide edge roughness caused significant waveguide-to-waveguide coupling between single and ensemble waveguides, cross-talk would be similar for ion (i) and (ii). As shown in Fig. 4a (dark blue), cross-talk was proximity dependent and therefore not significantly caused by the input couplers or waveguide-to-waveguide coupling on chip, but rather was dominated by scattered light from the output coupler. Another experiment mehta:2020 measured lower cross-talk (-60 dB) with gratings that emitted 729 nm light in the -1 order mode, which suppresses higher order modes. Given our lithography tolerances and a shorter 435 nm wavelength, we were limited to designing gratings that emit in the +1 order mode, and subsequently generated higher order modes that resulted in greater optical crosstalk. In addition, shorter wavelengths are more sensitive to fabrication imperfections that can lead to scattering.
Fig. 4b shows the data for the opposite case, where the two ion ensemble was addressed. The single ion measured a relative cross-talk intensity of $I_{c}/I_{0}=0.0036(2)$. The two measurements of cross-talk have similar intensities which suggests that the source of the cross-talk has little directional dependence.
V TOWARD A FULLY-INTEGRATED SYSTEM
Our trap was equipped with SPADs at each trap site for integrated single-site detection. The SPADs integrated in this trap were the same as previously designed setzer:2021 , however an aluminum mesh was added to shield the ion from the electric field of the SPAD (see Methods). This change allowed for trapping ${}^{171}\textrm{Yb}^{+}$ directly above the SPAD while it was active. We operated one SPAD independently and measured a count rate of over $7$ kcps for one quadrant of the SPAD, with dark counts below $1$ kcps. These results are better than the previous SPAD performance reported for ${}^{174}\textrm{Yb}^{+}$ setzer:2021 . Unfortunately, scattered light originating from the waveguides and output couplers propagated through the various layers of the device, saturating the SPADs and prohibiting integrated state-readout. In the future, features such as light baffles could mitigate this scatter and enable simultaneous fully integrated addressing and state-readout.
VI Conclusions
In this work, we demonstrated the individual addressing of trapped-ions at multiple sites on a surface trap using MMI waveguide splitters and fully integrated multi-color waveguides for cooling, state preparation, coherent operations and detection. This combination of integrated waveguide delivery and multi-ion control with single channel optical inputs shows a path towards scaling trapped-ion systems like quantum computers, where potentially millions of qubits reiher:2017 will be needed for a truly useful general-purpose quantum computer.
A more near-term application of this work is to create a robust and portable optical ion clock by replacing the optomechanics in lab-scale systems with integrated photonics like lasers-on-chip Elshaari:2020Npho , modulators hogle:2023 , detectors setzer:2021 ; Reens:2022PRL , and the waveguides and splitters described here. They can also be used to improve the clock accuracy by supporting multiple ensembles of ions that can be concurrently probed with different interrogation times, and by eliminating deadtime Schioppo:2017NPho with ensembles that have non-overlapping detection windows. Clock performance can be further enhanced by probing systematics Roseband:2013arx ; Kim:2022NP while simultaneously operating the clock.
Acknowledgements.
We thank the members of Sandia’s Microsystems and Engineering Sciences Application (MESA) facility for their fabrication expertise and for helpful comments on the manuscript. This work was supported by the Defense Advanced Research Projects Activity (DARPA). Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under Contract No. DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
Author Contributions.
J.K., W.J.S., D.L.S. and H.J.M. conceived the experiment. J.K. and W.J.S. took the measurement and analyzed data. M.G. designed the integrated optics while also performing waveguide tests and simulations with N.K. . J.V.D.W. developed multi-ion solution. R.L. setup the Doppler laser system. J.K., W.J.S., D.S. and H.J.M. wrote the original manuscript. All the authors reviewed and edited the manuscript. H.J.M. supervised the project.